Of i have a 75% in a class and get a 90% on a test, whats my new grade
Answer by stanbon(57290)

If I have a 75% in a class and get a 90% on a test, whats my new grade
---
1st test: 75 points out of 100
2nd test: 90 points out of 100
----
Average: (75+90)/200 = 165/200 = 82.5%
========================================
Cheers,
Stan H.
==============